• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name saw

    Log entry time 06:32:01 on March 24, 2005

    Entry number 142278

    keyword=Singles rates, beam quality

    After giving up on the compton, MCC gave us 10 uA. We noted that the rates in the neutron bars at 10uA were comparable to earlier rates at 40uA. They tried to tune without much effect. While they were trying to tune, Doug talked to MCC and found out that the quad at 1C20 was not restored to what it was two days ago.

    Requested that the 1C20 quad setting be restored. This cut rates about half, (still twice the earlier rates).

    So for now, are requesting 20uA to take a little bit of data with the new higher pm kinematics so we can see what the (e,e'p) trigger rate is.